OpenStudy (anonymous):

6 neighbors share 4 pies how much does each get

OpenStudy (cgreenwade2000):

What is 6 divided by 4

OpenStudy (anonymous):

6/4=3/2 simplify 3/2=1.5 one and a half pie each